# Birational Geometry of Special Quotient Foliations and Chazy’s Equations

### v1

@@ 4.1. Chazy IV and $\mathcal{G}_{\mathrm{IV}}$
Let us analyze the invariant curves and the singularities of $\mathcal{G}_{\mathrm{IV}}$. For calculations in the smooth chart $[1:x:y]$ of $\mathbb{P}(1,2,3)$, biholomorphic to $\mathbb{C}^{2}$, one can simply restrict (22) to $\{x=1\}$. For calculations in the singular charts of $\mathbb{P}(1,2,3)$, one can resort to the formulas for the desingularizations of $A_{1}$ and $A_{2}$ discussed in Section 2.2. Both $B$ and $C$ give invariant curves for $\mathcal{G}_{\mathrm{IV}}$; we will denote them by the same symbols. The foliation $\mathcal{G}_{\mathrm{IV}}$ is regular at both $p_{1}$ and $p_{2}$, in the sense that, at each one of these points, it is the quotient of a regular foliation. In the smooth part of $\mathbb{P}(1,2,3)$, $\mathcal{G}_{\mathrm{IV}}$ has three non-degenerate singularities:
  • $q_{1}=[1:0:0]$, with eigenvalues $-\omega:1$;
@@
  • $q_{3}=[2:2:4]$, a saddle with eigenvalues $-3:2$.
The curve $C$ has two transverse smooth branches at $q_{1}$. The curves $B$ and $C$ are smooth and tangent at $q_{2}$, where they are tangent to the eigenvalue $1$. The curve $B$ passes also through $p_{1}$ and $q_{3}$. Since there are two smooth invariant curves through $q_{2}$ tangent to each other, the foliation is linearizable in a neighborhood of this singularity, and the curves have a contact of order $3$.
After desingularizing $p_{1}$ and $p_{2}$, and resolving the singularity of the foliation at $q_{2}$, we find three chains of two rational curves of self-intersection $-2$ each:
@@
(see the bottom left and the top of Figure 5). The contraction of each one of these chains gives a singularity of type $A_{2}$. These give the three singularities of the quotient model for $\mathcal{F}_{3}$. The link comes from the transform of $C$.
@@
4.1.1. Birational equivalence between $\mathcal{G}_{\mathrm{IV}}$ and $\mathcal{F}_{3}$
Under the birational map (14), $\mathcal{G}_{\mathrm{IV}}$ induces the foliation of degree two on $\mathbb{P}^{2}$ given by (1), giving an alternative proof of Theorem 1. We remark that there is a full one-dimensional system (a pencil) of degree-two foliations leaving invariant the nodal cubic and the two inflectional tangents of Theorem 1, so, unlike the situation in Theorems 2 and Theorem 3, the foliation is not determined by its invariant algebraic curves.
@@ 4.1.1. Birational equivalence between $\mathcal{G}_{\mathrm{IV}}$ and $\mathcal{F}_{3}$
Let us describe the singularities of the foliation $\mathcal{H}_{3}$ on $\mathbb{P}^{2}$ given by (2). The singularities of $\mathcal{G}_{\mathrm{IV}}$ away from $\{x=0\}$ coming from $q_{1}$, $q_{2}$ and $q_{3}$, are respectively placed at $(1:0:0)$, $(1:1:2)$ and $(2:1:1)$, and admit the same local description. On the line $x=0$, we have the singular points $(0:1:3)$, a saddle with eigenvalues $-1:2$, and the point $(0:0:1)$, a nilpotent singularity with multiplicity three.
The composition of the resolution $\pi:S\to\mathbb{P}(1,2,3)$ with the map $j$ in (14) gives a resolution of this foliation of $\mathbb{P}^{2}$. See the right-hand side of Figure 5.
4.1.2. Relation with Pereira’s model
In [Per05, Sect. 5], Pereira gave a projective model for Brunella’s very special foliation $\mathcal{F}_{3}$. It is the degree-two foliation given by the vanishing of
@@ 4.1.2. Relation with Pereira’s model
[…]
@@
4.2. Chazy V and $\mathcal{G}_{\mathrm{V}}$
The associated vector field $W$ on $\mathbb{C}^{3}$ has the quasihomogeneous first integral of degree four $B=x^{4}-4x^{2}y+2zx-y^{2}$, and the invariant surface given by the quasihomogeneous polynomial of degree six
@@ 4.2. Chazy V and $\mathcal{G}_{\mathrm{V}}$
Let $\Sigma=B^{-1}(1)$. It has a natural action of $\mathbb{Z}_{4}$ preserving the foliation induced by $W$, given by the restriction of (7) to the group of fourth roots of unity.
@@
$(\mathrm{i}-1)\left(\mathrm{i}X\frac{\partial}{\partial X}+Y\frac{\partial}{\partial Y}\right),$
and which is tangent to the foliation $\mathcal{E}_{4}$ of Eq. (19) in Section 3.2.
Consider the map $\Phi_{4}:\mathbb{P}^{1}\times\mathbb{P}^{1}\dashrightarrow\Sigma$ given by $\Phi_{4}(X:1,Y:1)=(f,D_{4}f,D_{4}^{2}f),$ for
$f(X,Y)=\frac{(X-1)(Y-1)}{(XY+\mathrm{i}X-\mathrm{i}Y-1)}.$
It is equivariant with respect to the previously described action of $\mathbb{Z}_{4}$ on $\Sigma$, and to its action on $\mathbb{P}^{1}\times\mathbb{P}^{1}$ given by the transformation $T_{4}$ of Eq. (20). The map $\Phi_{4}$ is a birational one, and maps $D_{4}$ to the restriction of $W$ to $\Sigma$. It induces a birational equivalence between $\mathcal{G}_{\mathrm{V}}$, the foliation induced by $W$ on $\mathbb{P}(1,2,3)$, and $\mathcal{F}_{4}$ (the above formula for $f$ corrects the one given in [Gui12, p. 72]).
Let us now describe the singularities and invariant curves of $\mathcal{G}_{\mathrm{V}}$. There are invariant curves coming from $B$ and $C$, which we will denote be the same symbols. The foliation is regular at $p_{1}$ and $p_{2}$; away from these, its singularities are
@@
  • $q_{3}=[3:3:6]=$, a saddle with eigenvalues $-4:3$.
The curve $C$ has a node at $q_{1}$. At $q_{2}$, $B$ and $C$ are tangent, and tangent to the eigenvalue $1$. In particular, $q_{2}$ is linearizable, and the curves have a contact of order $4$. The curve $B$ passes also through $p_{2}$ and $q_{3}$.
After desingularizing $p_{2}$ and resolving $q_{2}$, we obtain two chains of rational curves of self intersection $-2$ of length three:
@@
(see the left-hand side of Figure 6). Upon contraction, they form the two singularities of type $A_{3}$ which, together with $p_{1}$, which is of type $A_{1}$, give the three singularities in the quotient model for $\mathcal{F}_{4}$.
@@
The birational involution coming from (21) reads
(27) $[x:y:z]\mapsto[x^{3}-3xy+z:-x^{4}y+6x^{2}y^{2}-4xyz-y^{3}+z^{2}:\\ -x^{6}z+8x^{5}y^{2}-x^{4}yz-32x^{3}y^{3}+29x^{2}y^{2}z+8xy^{4}-12xyz^{2}-3y^{3}z+2z^{3}].$
It is associated to the birational involution of $\mathbb{C}^{3}$ behind Theorem 7, which can be established by a direct computation.
4.2.1. Birational equivalence with $\mathcal{F}_{4}$
Under the birational map (14), $\mathcal{G}_{\mathrm{V}}$ is mapped to the degree-two foliation $\mathcal{H}_{4}$ of $\mathbb{P}^{2}$ given by (2). This establishes Theorem 3. The fact that the foliation is the unique one of degree two on $\mathbb{P}^{2}$ tangent to the cubic, conic and line follows from the fact that the tangency divisor of a pair of degree-two foliations on $\mathbb{P}^{2}$ has degree five.
Let us now study the singularities of the foliation. The previously described singular points $q_{1}$, $q_{2}$ and $q_{3}$ are, respectively, $(1:0:0)$, $(1:1:2)$ and $(9:3:2)$. On the line $x=0$, we have two singular points: $(0:1:4)$, a saddle with eigenvalues $-1:2$, and $(0:0:1)$, a nilpotent singularity with multiplicity three. The desingularization of $\mathcal{H}_{4}$ is given by the composition of the resolution $\pi:S\to\mathbb{P}(1,2,3)$ with the map $j$ in (14). See the right-hand side of Figure 6.
By conjugating the birational involution (27) via the birational map (14), we obtain:
@@ 4.2.1. Birational equivalence with $\mathcal{F}_{4}$
The Jacobian of the latter is
$4(x-y)^{4}(-3y+z+x)^{3}(x^{2}-4yx+2zx-y^{2}),$
with the line $x-3y+z=0$ being tangent to the conic at $(1:1:2)$; the lines $y-x=0$ and $x-3y+z=0$ intersecting at $(1:1:2)$; and the line $y-x=0$ intersecting the conic at $(0:0:1)$ and $(1:1:2)$. Its homaloidal system is formed by plane quartics having ordinary triple points at $(1:1:2)$, and being smooth at $(0:0:1)$. One of the local branches at $(1:1:2)$ defines a contact direction of the elements of the system: four blow-ups are need to separate these elements, one at $(1:1:2)$ and three along the directions given by the contact branch. At $(0:0:1)$, three blow ups are need to separate the smooth curves. In the sense of [NTN20], it is a de Jonquières map of type $(4;3,1^{3};1^{3})$, i.e., of type 32.1 in Table 5.1, p. 96.
4.3. Chazy VI and $\mathcal{G}_{\mathrm{VI}}$
The associated vector field $W$ has the first integral
@@ 4.3. Chazy VI and $\mathcal{G}_{\mathrm{VI}}$
of degree six, and the invariant hypersurface given by the quasihomogeneous polynomial of degree six
@@
These induce invariant curves for $\mathcal{G}_{\mathrm{VI}}$, the foliation induced by $W$ on $\mathbb{P}(1,2,3)$. The foliation $\mathcal{G}_{\mathrm{VI}}$ is regular at both $p_{1}$ and $p_{2}$; away from these, its singularities are

### v2

@@ 4.1.3. Description of $\mathcal{G}_{\mathrm{IV}}$
Let us analyze the invariant curves and the singularities of $\mathcal{G}_{\mathrm{IV}}$. For calculations in the smooth chart $[1:x:y]$ of $\mathbb{P}(1,2,3)$, biholomorphic to $\mathbb{C}^{2}$, one can simply restrict (22) to $\{x=1\}$. For calculations in the singular charts of $\mathbb{P}(1,2,3)$, one can resort to the formulas for the desingularizations of $A_{1}$ and $A_{2}$ discussed in Section 2.2. Both $B$ and $C$ give invariant curves for $\mathcal{G}_{\mathrm{IV}}$, which we will denote by the same symbols. On each one of the singular points $p_{1}$ and $p_{2}$ of $\mathbb{P}(1,2,3)$, the foliation $\mathcal{G}_{\mathrm{IV}}$ is regular, in the sense that, at each one of these points, it is the quotient of a regular foliation. In the smooth part of $\mathbb{P}(1,2,3)$, $\mathcal{G}_{\mathrm{IV}}$ has three non-degenerate singularities:
  • $q_{1}=[1:0:0]$, with eigenvalues $\omega:1$;
@@
  • $q_{3}=[2:2:4]$, a saddle with eigenvalues 3:2-3:2.
The curves $B$ and $C$ are smooth and tangent at $q_{2}$; since there are two smooth invariant curves through this point tangent to each other, the foliation is linearizable in a neighborhood of it, and the curves have a contact of order $3$. The curve $C$ passes also through $q_{1}$, where it has two transverse smooth branches; the curve $B$ passes also through $p_{1}$ and $q_{3}$.
After desingularizing $p_{1}$ and $p_{2}$, and resolving the singularity of the foliation at $q_{2}$, we find three chains of two rational curves of self-intersection 2-2 each:
@@
(see the bottom-left and the top of Figure 5). The contraction of each one of these chains gives a singularity of type $A_{2}$. These give the three singularities of the quotient model for $\mathcal{F}_{3}$. The link comes from the transform of $C$.
@@
4.1.4. The plane foliation $\mathcal{H}_{3}$
The birational map (14) transforms the foliation $\mathcal{G}_{\mathrm{IV}}$ into the foliation $\mathcal{H}_{3}$ of degree two on $\mathbb{P}^{2}$ given by (1). This gives an alternative proof of Theorem 1.
Remark 14.
There is a full one-dimensional system (a pencil) of degree-two foliations leaving invariant the nodal cubic and the two inflectional tangents of Theorem 1, so, unlike the situation in Theorems 2 and 3, the foliation is not determined by its invariant algebraic curves.
The invariant curve $x=0$ of $\mathcal{H}_{3}$ is produced by the birational modification (14) from $\mathbb{P}(1,2,3)$ to $\mathbb{P}^{2}$. The invariant curves $B$ and $C$ for $\mathcal{G}_{\mathrm{IV}}$ produce, for $\mathcal{H}_{3}$, the other invariant line and the invariant cubic in the statement Theorem 1.
Let us describe the singularities of the foliation $\mathcal{H}_{3}$ on $\mathbb{P}^{2}$ given by (1). The singularities of $\mathcal{G}_{\mathrm{IV}}$ away from $x=0$, $q_{1}$, $q_{2}$ and $q_{3}$, give singularities for $\mathcal{H}_{3}$, placed at $(1:0:0)$, $(1:1:2)$ and $(2:1:1)$, which admit the same local descriptions. On the line $x=0$ of $\mathbb{P}^{2}$, the singularities of $\mathcal{H}_{3}$ are $(0:1:3)$, a saddle with eigenvalues 1:2-1:2, and the point $(0:0:1)$, a nilpotent singularity with multiplicity three.
The composition of the resolution $\varpi:S\to\mathbb{P}(1,2,3)$ of Section 2.3 with the map $j$ of Eq. (14) gives a resolution of $\mathcal{H}_{3}$; see the right-hand side of Figure 5.
@@ 4.1.4. The plane foliation $\mathcal{H}_{3}$
4.1.5. Relation of $\mathcal{H}_{3}$ with Pereira’s model for $\mathcal{F}_{3}$
In [Per05, Sect. 5], Pereira gave a projective model for Brunella’s very special foliation $\mathcal{F}_{3}$. It is the degree-two foliation given by
@@ 4.1.5. Relation of $\mathcal{H}_{3}$ with Pereira’s model for $\mathcal{F}_{3}$
The foliation (1) is linearly equivalent to Pereira’s model for $\mathcal{F}_{3}$ (28), via the linear map (x:y:z)=(X:XY:2X3Y+Z)(x:y:z)=(X:X-Y:2X-3Y+Z), with inverse (X:Y:Z)=(x:xy:x3y+z)(X:Y:Z)=(x:x-y:x-3y+z).
@@
4.2. The Chazy V equation and Santos’s foliation $\mathcal{F}_{4}$
In this section we will establish the part of Theorem 4 concerning the birational equivalence between $\mathcal{F}_{4}$ and the foliation on $\mathbb{P}(1,2,3)$ given by the Chazy V equation. Theorems 2 and 7 are consequences of this equivalence, and will be established here as well.
4.2.1. Chazy V and $\mathcal{F}_{4}$
The Chazy V equation (5) is given by the quasihomogeneous vector field on $\mathbb{C}^{3}$
$W=y\frac{\partial}{\partial x}+z\frac{\partial}{\partial y}+(2xz+4y^{2}-2x^{2}y)\frac{\partial}{\partial z}.$
The latter induces a foliation on $\mathbb{P}(1,2,3)$; it will be denoted by $\mathcal{G}_{\mathrm{V}}$. The discussion that follows is based on [Gui12]. The vector field has the quasihomogeneous first integral of degree four $B=x^{4}-4x^{2}y+2zx-y^{2}$, and the invariant surface defined by the quasihomogeneous polynomial of degree six
@@ 4.2.1. Chazy V and $\mathcal{F}_{4}$
Let $\Sigma=B^{-1}(1)$. The foliation on $\Sigma$ induced by $W$ will be denoted by $\widetilde{\mathcal{G}}_{\mathrm{V}}$. There is a natural action of $\mathbb{Z}_{4}$ on $\Sigma$, given by the restriction of (7) to the group of fourth roots of unity, generated by $(x,y,z)\mapsto(\mathrm{i}x,-y,-\mathrm{i}z)$, which preserves $\widetilde{\mathcal{G}}_{\mathrm{V}}$.
@@
$(\mathrm{i}-1)\left(\mathrm{i}X\frac{\partial}{\partial X}+Y\frac{\partial}{\partial Y}\right).$
It is tangent to the foliation $\mathcal{E}_{4}$ of Eq. (19) in Section 3.2. As we there discussed, the action of $\mathbb{Z}_{4}$ on $\mathbb{P}^{1}\times\mathbb{P}^{1}$ given by the transformation $T_{4}$ of Eq. (20) produces, after quotient, a surface endowed with a foliation, Santos’s foliation $\mathcal{F}_{4}$.
Let
$f(X,Y)=\frac{(X-1)(Y-1)}{(XY+\mathrm{i}X-\mathrm{i}Y-1)},$
and consider the birational map $\Phi_{4}:\mathbb{P}^{1}\times\mathbb{P}^{1}\dashrightarrow\Sigma$ given by $\Phi_{4}(X:1,Y:1)=(f,D_{4}f,D_{4}^{2}f)$. The map $\Phi_{4}$ maps $D_{4}$ to the restriction of $W$ to $\Sigma$, thus mapping $\mathcal{E}_{4}$ to $\widetilde{\mathcal{G}}_{\mathrm{V}}$. It is equivariant with respect to the previously described actions of $\mathbb{Z}_{4}$ on $\Sigma$ and on $\mathbb{P}^{1}\times\mathbb{P}^{1}$. Thus, $\Phi_{4}$ induces the birational equivalence between $\mathcal{G}_{\mathrm{V}}$ and Santos’s foliation $\mathcal{F}_{4}$ stated in Theorem 4. (The above formula for $f$ corrects the one given in [Gui12, p. 72]; the one given there for the inverse of $\Phi_{4}$ is correct.)
4.2.2. Birational symmetries of Chazy V and $\mathcal{G}_{\mathrm{V}}$
By conjugating the involution (21) by $\Phi_{4}$ and its inverse, we obtain a birational involution of $\Sigma$, which produces, on its turn, the birational involution of $\mathbb{P}(1,2,3)$
(29) $x:y:z]\mapsto[x^{3}-3xy+z:-x^{4}y+6x^{2}y^{2}-4xyz-y^{3}+z^{2}:\\ -x^{6}z+8x^{5}y^{2}-x^{4}yz-32x^{3}y^{3}+29x^{2}y^{2}z+8xy^{4}-12xyz^{2}-3y^{3}z+2z^{3}].$
By a procedure in all ways similar to that of the previous section, this involution may be promoted to a birational involution of $\mathbb{C}^{3}$ preserving the vector field giving the Chazy V equation. This is the content Theorem 7, which can be established by a direct computation, and which we omit.
4.2.3. Description of $\mathcal{G}_{\mathrm{V}}$
The curves on $\mathbb{P}(1,2,3)$ defined by $B$ and $C$ are invariant by $\mathcal{G}_{\mathrm{V}}$; we will denote them by the same symbols. The foliation is regular at the singular points $p_{1}$ and $p_{2}$ of $\mathbb{P}(1,2,3)$; away from these, its singularities are:
@@ 4.2.3. Description of $\mathcal{G}_{\mathrm{V}}$
  • $q_{3}=[3:3:6]$, a saddle with eigenvalues 4:3-4:3.
The curves $B$ and $C$ pass through the point $q_{2}$, at which they are smooth and tangent. In particular, $q_{2}$ is a linearizable singularity of the foliation, and the curves have a contact of order $4$. The curve $C$ passes also through $q_{1}$, where it has a node, and $B$ passes also through $p_{2}$ and $q_{3}$.
After desingularizing $p_{2}$, and resolving the singularity of the foliation at $q_{2}$, we obtain two chains of length three of rational curves of self intersection 2-2:
@@
(see the left-hand side of Figure 6). Upon contraction, they form the two singularities of type $A_{3}$ which, together with $p_{1}$ (which is of type $A_{1}$), give the three singularities in the quotient model for $\mathcal{F}_{4}$ described in Section 3.2.
@@
4.2.4. Birational equivalence of $\mathcal{F}_{4}$ and $\mathcal{H}_{4}$
Under the birational map (14), $\mathcal{G}_{\mathrm{V}}$ is mapped to the degree-two foliation $\mathcal{H}_{4}$ of $\mathbb{P}^{2}$ given by (2). The invariant line $x=0$ of $\mathcal{H}_{4}$ is produced by the birational map (14); the other invariant curves in the statement of Theorem 2 are the strict transforms of $B$ and $C$ under (14). The fact that $\mathcal{H}_{4}$ is the only foliation of degree two on $\mathbb{P}^{2}$ tangent to the cubic, conic and line, follows from the fact that the tangency divisor of a pair of degree-two foliations on $\mathbb{P}^{2}$ has degree five. This establishes Theorem 2.
Let us describe the singularities of $\mathcal{H}_{4}$. The previously described singular points $q_{1}$, $q_{2}$ and $q_{3}$ of $\mathcal{G}_{\mathrm{V}}$ become, respectively, the singular points $(1:0:0)$, $(1:1:2)$ and $(9:3:2)$ of $\mathcal{H}_{4}$. On the line $x=0$, $\mathcal{H}_{4}$ has two further singular points: $(0:1:4)$, a saddle with eigenvalues 1:2-1:2, and $(0:0:1)$, a nilpotent singularity with multiplicity three. The desingularization of $\mathcal{H}_{4}$ is given by the composition of the resolution $\varpi:S\to\mathbb{P}(1,2,3)$ with the map $j$ in (14). See the right-hand side of Figure 6.
By conjugating the birational involution (29) by the birational map (14), we obtain:
@@ 4.2.4. Birational equivalence of $\mathcal{F}_{4}$ and $\mathcal{H}_{4}$
The Jacobian of a birational map of $\mathbb{P}^{2}$ is the determinant of the Jacobian matrix of its lift to $\mathbb{C}^{3}$. Its vanishing gives the projective curves contracted to points by the birational map (cf. Prop. 3.5.3 of [ACFLl21]). For the degree four map above, the Jacobian is
$4(x-y)^{4}(x-3y+z)^{3}(x^{2}-4yx+2zx-y^{2}),$
with the line x3y+z=0x-3y+z=0 being tangent to the conic at $(1:1:2)$; the lines yx=0y-x=0 and x3y+z=0x-3y+z=0 intersecting at $(1:1:2)$; and the line yx=0y-x=0 intersecting the conic at $(0:0:1)$ and $(1:1:2)$.
In general, a birational map of $\mathbb{P}^{2}$ determines a homaloidal system (a net of rational curves that map to the straight lines of $\mathbb{P}^{2}$ under it). When the homaloidal system of a degree $d$ birational map has base-points of multiplicity one except for one base-point $O$ with multiplicity d1d-1, the map is called a de Jonquières one, and can be characterized as a birational map preserving the pencil of straight lines by $O$.
The homaloidal system of the above degree-four map is formed by quartics having ordinary triple points at $(1:1:2)$, and that are smooth and tangent at $(0:0:1)$. One of the local branches at $(1:1:2)$ defines a contact direction of the elements of the system: four blow-ups are needed to separate these elements, one at $(1:1:2)$ and three along the directions given by the contact branch. At $(0:0:1)$, three blow ups are needed to separate the curves. Thus, it is a de Jonquières map of type 32.1 in Table 5.1, p. 96, of [NTN20].
4.3. The Chazy VI equation and Santos’s foliation $\mathcal{F}_{6}$
In this section we will establish the birational equivalence between Santos’s foliation $\mathcal{F}_{6}$ and the foliation on $\mathbb{P}(1,2,3)$ induced by the Chazy VI equation stated in Theorem 4. As a consequence, we will establish Theorem 3. Theorem 8 will follow from it and from Remark 13.
4.3.1. Chazy VI and $\mathcal{F}_{6}$
The Chazy VI equation (6) is given by the quasihomogeneous vector field on $\mathbb{C}^{3}$
$W=y\frac{\partial}{\partial x}+z\frac{\partial}{\partial y}+(xz+5y^{2}-x^{2}y)\frac{\partial}{\partial z}.$
It induces a foliation on $\mathbb{P}(1,2,3)$, that we will denote by $\mathcal{G}_{\mathrm{VI}}$. The discussion that follows is based on [Gui12]. The vector field has the first integral of degree six
@@ 4.3.1. Chazy VI and $\mathcal{F}_{6}$
and the invariant hypersurface given by the quasihomogeneous polynomial of degree six
@@
For $\Sigma=B^{-1}(1)$, we have a birational map $\Phi_{6}:\mathrm{Bl}_{3}(\mathbb{P}^{2})\to\Sigma$ [Gui12, p. 73], equivariant with respect to the actions of $\mathbb{Z}_{6}$ given the action of sixth roots of unity on $\Sigma$ via (7), and the action of $T_{6}$ on $\mathrm{Bl}_{3}(\mathbb{P}^{2})$ described in Section 3.3. It maps $\mathcal{E}_{6}$ to the foliation induced by $W$, and induces an identification of $\mathcal{G}_{\mathrm{VI}}$ with $\mathcal{F}_{6}$.
4.3.2. Relation between $\mathcal{G}_{\mathrm{IV}}$ and $\mathcal{G}_{\mathrm{VI}}$
The relation between $\mathcal{F}_{4}$ and $\mathcal{F}_{6}$ discussed in Remark 13 has a counterpart for for the Chazy IV and VI equations and the foliations they induce on $\mathbb{P}(1,2,3)$. From the explicit expressions for $\Phi_{6}$ and the inverse of $\Phi_{3}$ in [Gui12, p. 71], we may realize an explicit two-to-one map from $\mathbb{P}(1,2,3)$ to itself, mapping $\mathcal{G}_{\mathrm{IV}}$ to $\mathcal{G}_{\mathrm{VI}}$:
$x:y:z]\mapsto[xz-2y^{2}:3x^{4}y^{2}-3x^{3}yz-9x^{2}y^{3}+x^{2}z^{2}+8xy^{2}z+4y^{4}-3yz^{2}:\\ 9x^{6}y^{3}-12x^{5}y^{2}z-45x^{4}y^{4}+6x^{4}yz^{2}+63x^{3}y^{3}z-x^{3}z^{3}+\\ +54x^{2}y^{5}-42x^{2}y^{2}z^{2}-48xy^{4}z+15xyz^{3}-16y^{6}+18y^{3}z^{2}-3z^{4}].$
In a way similar to that of Theorems 6 and 7, this rational map can be promoted to a rational map of $\mathbb{C}^{3}$ onto itself mapping the vector field of the Chazy IV equation to that of the Chazy VI one. This is the statement of Theorem 8. It can be established by a direct calculation, and we omit its proof.
4.3.3. Description of $\mathcal{G}_{\mathrm{VI}}$
We have invariant curves for $\mathcal{G}_{\mathrm{VI}}$ given by $B$ and $C$. The foliation $\mathcal{G}_{\mathrm{VI}}$ is regular at the singular points $p_{1}$ and $p_{2}$ of $\mathbb{P}(1,2,3)$; away from these, its singularities are